Problem 6-03A a, b1-b2, c (Part Level Submission) (Video)

Oriole Company had a beginning inventory on January 1 of 323 units of Product 4-18-15 at a cost of $20 per unit. During the year, the following purchases were made.

Mar. 15 860 units at $23 Sept. 4 753 units at $26
July 20 538 units at $24 Dec. 2 215 units at $29

2,150 units were sold. Oriole Company uses a periodic inventory system.

Determine the cost of goods available for sale.

The cost of goods available for sale $

Entry field with correct answer64,965

Calculate average cost per unit. (Round answer to 2 decimal places, e.g. 2.23.)

Average cost per unit $24.16

Entry field with incorrect answer

Determine (1) the ending inventory, and (2) the cost of goods sold under each of the assumed cost flow methods (FIFO, LIFO, and average-cost). Prove the accuracy of the cost of goods sold under the FIFO and LIFO methods. (Round answers to 0 decimal places, e.g. 1,250.)

FIFO

LIFO

AVERAGE-COST

The ending inventory $

$

$

The cost of goods sold $

$

$
Answer #1
Cost of Goods available for sale
Units Cost/unit COG for sale
Beginning Inventory 323 20 6460
Purchases:
Mar, 15 860 23 19780
july 20, 538 24 12912
Sept, 4 753 26 19578
Dec, 2 215 29 6235
TOTAL 2689 64965
Average cost per unit = Cost of goods available for sale /Total units = 64965/2689 = 24.16
Requirement ‘b’
FIFO Cost of Goods available for sale Cost of Goods Sold Ending Inventory
A B Cost/unit D E F H H I J
Units Cost/unit COG for sale (A*B) Units sold Cost/unit COGS = (E*F) Units Cost/unit Ending inventory = (H*I)
Beginning Inventory 323 20 6460 323 20 6460 0 20 0
Purchases:
Mar, 15 860 23 19780 860 23 19780 0 23 0
july 20, 538 24 12912 407 24 9768 131 24 3144
Sept, 4 753 26 19578 26 0 753 26 19578
Dec, 2 215 29 6235 29 0 215 29 6235
TOTAL 2689 64965 1590 36008 1099 28957
LIFO Cost of Goods available for sale Cost of Goods Sold Ending Inventory
A B Cost/unit D E F H H I J
Units Cost/unit COG for sale (A*B) Units sold Cost/unit COGS = (E*F) Units Cost/unit Ending inventory = (H*I)
Beginning Inventory 323 20 6460 0 20 0 323 20 6460
Purchases:
Mar, 15 860 23 19780 84 23 1932 776 23 17848
july 20, 538 24 12912 538 24 12912 0 24 0
Sept, 4 753 26 19578 753 26 19578 0 26 0
Dec, 2 215 29 6235 215 29 6235 0 29 0
TOTAL 2689 64965 1590 40657 1099 24308
Average Method Cost of Goods available for sale
Units Cost/unit COG for sale (A*B)
Beginning Inventory 323 20 6460
Purchases:
Mar, 15 860 23 19780
july 20, 538 24 12912
Sept, 4 753 26 19578
Dec, 2 215 29 6235
TOTAL 2689 64965
Average cost per unit = Cost of goods available for sale /Total units = 64965/2689 = 24.16
               
Cost of goods sold = Units sold * Average cost per unit = 1590*24.16 =38414
Ending inventory = Balance units * Average cost per unit = 1099*24.16 =26551
FIFO LIFO Average Cost
Ending Inventory 28957 24308 26551
Cost of Goods Sold [ units] 36008 40657 38414
Total 64965 64965 64965